# A Study of Library Usage in Agent-Authored Pull Requests

### v1

@@ Abstract.
Coding agents are becoming increasingly capable of completing end-to-end software engineering workflows that previously required a human developer, including raising pull requests (PRs) to propose their changes. However, we still know little about how these agents use libraries when generating code, a core part of real-world software development. To fill this gap, we study 26,760 agent-authored PRs from the AIDev dataset to examine three questions: how often do agents import libraries, how often do they introduce new dependencies (and with what versioning), and which specific libraries do they choose? We find that agents often import libraries (29.5% of PRs) but rarely add new dependencies (1.3% of PRs); and when they do, they follow strong versioning practices (75.0% specify a version), an improvement on direct LLM usage where versions are rarely mentioned. Generally, agents draw from a surprisingly diverse set of external libraries, contrasting with the limited “library preferences” seen in prior non-agentic LLM studies. Our results offer an early empirical view into how AI coding agents interact with today’s software ecosystems.
@@ 1. Introduction
Large language models (LLMs) have rapidly reshaped the way developers write code, moving beyond single-line completions to generating whole functions or creating new projects (chenEvaluatingLargeLanguage2021; jiangSurveyLargeLanguage2024). More recently, the field has seen a shift from human-centric code completion to agentic workflows in which LLMs plan, act, and call tools to solve more realistic software engineering tasks (liuLargeLanguageModelBased2025). These agentic systems—now widely available in production tools (shrivastavaComparativeFeaturesetAnalysis2025)—can complete end-to-end workflows that previously required a human developer. Crucially, these workflows include raising pull requests (PRs), allowing agents to propose substantial code changes rather than the narrow dependency updates handled by earlier bots (heAutomatingDependencyUpdates2023a).
@@
For agents to contribute meaningfully to real projects, they must be able to work effectively with external libraries. This is central to modern software development (somervilleSoftwareEngineeringGlobal2016), where reusing established libraries improves efficiency and avoids reinventing well-tested functionality (xuWhyReinventingWheels2020). However, prior work shows that LLMs in non-agentic settings often struggle here: hallucinating package names (twistLibraryHallucinationsLLMs2025), calling deprecated APIs (wangLLMsMeetLibrary2025), or omitting version constraints altogether (rajRoleLibraryVersions2024a). LLMs often rely on retrieval-based solutions to overcome these issues (jainMitigatingCodeLLM2024; wangExploraCoderAdvancingCode2025), but agents have the autonomy to bypass this additional infrastructure. What remains unclear is whether agentic workflows will actually use external libraries—and if they do, whether this autonomy helps mitigate existing problems or simply reproduces them at scale.
@@
To begin to fill this gap, we present the first focused empirical study of how AI coding agents use libraries in practice, by mining real agent-authored PRs from the AIDev dataset (liRiseAITeammates2025). Specifically, we analyse file diffs (both code and dependency manifests) from 26,760 PRs across four languages (TypeScript, Python, Go, and C#), to answer three research questions (RQs):
@@
RQ1 Library Usage How often do agents import libraries when generating code, and do they favour built-in or external libraries?
@@
RQ2 New Dependencies How often do agents introduce new dependencies into a repository, and do they specify versions?
@@
RQ3 Choosing Libraries Which specific libraries do agents choose, and what functional roles do they play?
@@
Our findings are concise but instructive. (1) Agents often import libraries (29.5% of PRs) but rarely add entirely new dependencies (1.3% of PRs). (2) Agentic workflows encourage far better version hygiene than traditional LLM prompting: agents specify versions for 75.0% of added dependencies, whereas LLMs would only mention versions in $\approx$9% library-related interactions (rajRoleLibraryVersions2024a). (3) Agents use a surprisingly diverse range of libraries, importing 3,988 distinct external packages across 7,888 PRs—a sharp contrast to the narrow “library preferences” previously observed in more self-contained LLM tasks (twist-arxiv-llm-code-bias).
Our contributions are as follows:
@@
Programming Language Code File Extensions Example Import Statements Dependency Manifests Example Dependencies TypeScript (typescriptDocumentationModules2025) .ts, .tsx, .js, .jsx, .mjs, .cjs import { useEffect } from ’react’; import { Class } from ’@angular/core’; package.json "lodash": $\sim$4.17.21", Python (python5ImportSystem2025) .py import numpy as np from django.db import models requirements.txt requests pandas<2.0 pyproject.toml "pytest" "numpy==2.79", Go (goPackagesGoProgramming2025) .go import "fmt" import ( mux "github.com/gorilla/mux" ) go.mod require github.com/pkg/errors v0.9.1 C# (microsoftUsingDirectiveImport2025) .cs using Newtonsoft.Json; using Serilog; {name}.csproj packages.config <PackageReference     Include="Serilog" Version="2.12.0" /> paket.dependencies nuget Newtonsoft.Json 13.0.1
@@ 2. Dataset
To conduct our study, we use the AIDev dataset (liRiseAITeammates2025), a large-scale corpus of almost a million pull requests (PRs) from over 116,000 real GitHub repositories, authored by five distinct LLM-based coding agents (Claude Code, Cursor, Devin, Copilot, OpenAI Codex). Specifically, we use the November 2025 snapshot (commit eee0408 (hao-liHaoliAIDevEee0408a277826d88fc0ca5fa07d2fc325c96af12025)).
For our analysis, we use the AIDev-Pop subset that is curated to repositories with over 100 GitHub stars and includes the full commit details; AIDev-Pop contains 33,596 PRs from 2,807 repositories. We use three tables: repository, containing metadata about each repository; pull_request, which provides high-level PR information; and pr_commit_details, which contains file-level diffs and commit content. The dataset provides diffs in the “combined diff format” (gitGitGitdiffDocumentation), allowing us to extract new content from each file by selecting lines that begin with whitespace followed by exactly one “+”.
We restrict our study to the four most common languages in the dataset—those with the largest number of repositories—to ensure both depth of analysis and sufficient statistical support. The chosen languages are TypeScript/JavaScript (combined due to nearly identical syntax and library ecosystems, and will be referred to as TypeScript throughout the paper), Python, Go and C#. We initially included Rust (fifth most popular language), but its PRs relied heavily on internal workspace crates that were indistinguishable from external imports, so we limited our study to the top four languages.
Ethical considerations
@@
Metric TypeScript Python Go C# All languages General statistics Total repositories 840 530 242 220 1,832 Total pull requests 7,480 7,190 10,107 1,983 26,760 Total PRs merged 4,428 (59.2%) 5,468 (76.1%) 8,361 (82.7%) 1,146 (57.8%) 19,403 (72.5%) RQ1 Library Usage PRs importing any library 2,993 (40.0%) 2,720 (37.8%) 1,273 (12.6%) 902 (45.5%) 7,888 (29.5%) PRs importing standard library 828 (11.1%) 2,383 (33.1%) 1,244 (12.3%) 743 (37.5%) 5,198 (19.4%) PRs importing external libraries 2,847 (38.1%) 1,914 (26.6%) 324 (3.2%) 738 (37.2%) 5,823 (21.8%) Avg. standard library imports per PR 0.26 1.28 0.63 0.49 0.69 Avg. external library imports per PR 1.87 0.73 0.08 0.73 0.80 Unique external libraries imported 2,495 948 314 231 3,988 PRs with library imports merged 1,595 (53.3%) 1,768 (65.0%) 918 (72.1%) 467 (51.8%) 4,748 (60.2%) RQ2 New Dependencies PRs adding new dependencies 9 (0.1%) 200 (2.8%) 16 (0.2%) 120 (6.1%) 345 (1.3%) Avg. new dependencies per PR 0.01 0.17 0.02 0.30 0.08 Total new dependencies 95 1,237 203 593 2,128 Dependencies specifying version 95 (100.0%) 1,020 (82.5%) 203 (100.0%) 278 (46.9%) 1,596 (75.0%) Unique new dependencies 85 552 188 294 1,119 PRs with new dependencies merged 6 (66.7%) 131 (65.5%) 5 (31.2%) 56 (46.7%) 198 (57.4%)
@@ 3.1. Method for RQ1 Library Usage
We begin by investigating how inclined agents are to make use of libraries in their code. We recognise library usage only when imports appear in agent-generated code, because imports that already existed in a file prior to the agent’s edit do not reflect the agent’s preferences; and with only partial file snapshots provided in diffs, downstream function calls are often ambiguous without the corresponding import statement. We use file extensions to identify code files for each programming language, as defined by GitHub Linguist (github-linguistLinguistLanguageSavant).
We implement language-specific regex parsers to automatically extract import statements from those files, taking care to distinguish relative imports (from within the project itself). All extraction is supported by unit tests to ensure reliability. Information on the file extensions we consider, along with example imports, is provided in Table 1. Each library is categorised as either standard (built-in, as defined in the language documentation) or external (installed via a third-party package manager). This allows us to measure how heavily agents rely on external ecosystems versus native language capabilities.
@@ 3.2. Method for RQ2 New Dependencies
Next, we examine how often agents introduce new libraries into repositories. Here, we are particularly interested in the quality of these additions—specifically, whether agents are explicit with version constraints—because prior work shows that LLMs in non-agentic workflows rarely discuss versions (rajRoleLibraryVersions2024a), despite the known risks. For each language, we identify its corresponding dependency manifest files (for example, “requirements.txt” for Python, “package.json” for TypeScript) following the GitHub Docs (githubDependencyGraphSupported); we ignore lock files, which contain large numbers of second-hand dependencies not explicitly used within the repository.
We then implement manifest-specific regex parsers to automatically identify new dependencies—including any explicit version specifiers (for example, “==1.2” or “¿=1.0”)—again with unit tests to ensure reliability. The manifest files we consider, and example dependency definitions, are listed in Table 1. This enables us to quantify how often agents introduce completely new libraries and how reliably they do so.
@@ 3.3. Method for RQ3 Choosing Libraries
Finally, we examine the specific libraries agents choose and the functional roles they play. For each language, we take the set of libraries identified in RQ1 Library Usage and RQ2 New Dependencies and compute simple frequency statistics to surface the most commonly used or newly introduced libraries. We then perform manual analysis on the top ten libraries for each language and RQ to determine the areas that agents are most likely to use external libraries.
@@ 4. Findings & Discussion
We analyse a total of 26,760 agent-authored pull requests across 1,832 repositories and the four most common languages in the AIDev dataset. Here, we present and discuss our findings.
@@ 4.1. Findings for RQ1 Library Usage
For this RQ, we extract all libraries imported by agents and examine how often they rely on built-in standard libraries versus external third-party dependencies. The full breakdown is shown in Table 2.
Agents often import libraries when generating code, with 29.5% of PRs importing at least one library, and an average of 1.5 libraries imported across all PRs. However, the extent varies substantially by language: agents lean most heavily on libraries when working in TypeScript and C# (40.0% and 45.5% of PRs), while Go is a clear outlier, importing libraries in just 12.6% of PRs, suggesting that agents mostly modify the Go code without using additional tooling. This shows up again when looking at average imports per PR: both TypeScript and Python exceed two libraries per PR, whereas Go averages below one. Although reusing libraries is a core skill for professional developers, there is no established baseline for how often this should occur, making it difficult to assess whether these rates indicate overuse, underuse, or neither.
Agents rely strongly on the standard library in Python and C# (33.1% and 37.5% of PRs), but their use of external libraries in those languages is comparatively well balanced (26.6% and 37.2%). In TypeScript, the pattern is very different: agents overwhelmingly favour external packages (38.1% of PRs) over the standard library (11.1%), suggesting a strong inclination toward high-level tooling. Once again Go is the exception, using the standard library almost four times more often than external dependencies.
Finally, external library diversity is also high: agents collectively import 3,988 unique libraries. This contrasts sharply with earlier work showing that LLMs typically exhibit low library diversity (twist-arxiv-llm-code-bias). It suggests that giving agents rich project-level context nudges them toward much broader library use. While it may seem surprising that TypeScript alone contributes more than half of all unique libraries, this is entirely consistent with the ecosystem’s well-known fragmentation and reliance on micro-packages (kulaImpactMicroPackagesEmpirical2017).
RQ1 Summary
Agents often import libraries for TypeScript (40.0% of PRs), Python (37.8%) and C# (45.5%); and rarely for Go (12.6%). For TypeScript, agents rely on external libraries over built-in, for Python and C# we observe more balanced library usage, and for Go the standard library is highly preferred.
@@
Metric TypeScript Python Go C# Top 10 external library imports react (1013) vitest (483) zod (313) next (303) lucide-react (261) @testing-library/react (136) @tanstack/react-query (111) lodash (107) drizzle-orm (97) chai (92) pytest (812) pydantic (279) crewai (264) requests (125) numpy (125) fastapi (114) openai (65) pandas (60) httpx (53) yaml (52) github.com/stretchr/testify (108) github.com/smacker/go-tree-sitter (21) github.com/spf13/cobra (20) github.com/cometbft/cometbft (13) github.com/microsoft/typescript-go (13) github.com/tliron/glsp (13) go.uber.org/zap (12) github.com/gofiber/fiber (11) k8s.io/apimachinery/pkg (10) github.com/onsi/gomega (9) Avalonia (275) Xunit (243) Dock (131) Aspire (55) Azure (50) NUnit (48) FluentAssertions (30) ReactiveUI (28) Newtonsoft (27) Svg (27) Top 10 new dependencies typescript (3) vite (3) husky (2) @biomejs/biome (2) zod (2) prettier (2) @types/node (2) three (2) @electron-forge/plugin-auto-unpack-natives (1) eslint-plugin-react-compiler (1) requests (24) openai-agents (22) openai (21) numpy (21) pandas (18) pytest (14) python-dotenv (13) uvicorn (13) pydantic (12) aiohttp (12) github.com/golang/snappy (6) github.com/stretchr/testify (3) gopkg.in/yaml.v3 (3) github.com/klauspost/reedsolomon (3) github.com/gogo/protobuf (3) go.uber.org/atomic (2) github.com/smacker/go-tree-sitter (2) github.com/leonklingele/grouper (1) github.com/sashamelentyev/usestdlibvars (1) github.com/breml/bidichk (1) Microsoft.NET.Test.Sdk (26) xunit.runner.visualstudio (22) xunit (20) Moq (20) coverlet.collector (15) Microsoft.AspNetCore.Mvc.Testing (11) NUnit (9) NUnit3TestAdapter (9) Azure.Identity (7) Microsoft.Extensions.DependencyInjection (7)

### v2

@@ Abstract.
Coding agents are becoming increasingly capable of completing end-to-end software engineering workflows that previously required a human developer, including raising pull requests (PRs) to propose their changes. However, we still know little about how these agents use libraries when generating code, a core part of real-world software development. To fill this gap, we study 26,760 agent-authored PRs from the AIDev dataset to examine three questions: how often do agents import libraries, how often do they introduce new dependencies (and with what versioning), and which specific libraries do they choose? We find that agents often import libraries (29.5% of PRs) but rarely add new dependencies (1.3% of PRs); and when they do, they follow strong versioning practices (75.0% specify a version), an improvement on direct LLM usage where versions are rarely mentioned. Generally, agents draw from a surprisingly diverse set of external libraries, contrasting with the limited “library preferences” seen in prior non-agentic LLM studies. Our findings offer an early empirical view on how AI coding agents interact with today’s software ecosystems.
LLMs, AI coding agents, software libraries, dependency management, pull requests, empirical software engineering
journalyear: 2026copyright: ccconference: 23rd International Conference on Mining Software Repositories; April 13–14, 2026; Rio de Janeiro, Brazilbooktitle: 23rd International Conference on Mining Software Repositories (MSR ’26), April 13–14, 2026, Rio de Janeiro, Brazildoi: 10.1145/3793302.3793562isbn: 979-8-4007-2474-9/2026/04ccs: Software and its engineering Software libraries and repositoriesccs: Computing methodologies Intelligent agents
@@ 1. Introduction
Large language models (LLMs) have rapidly reshaped the way developers write code, moving beyond single-line completions to generating whole functions or creating new projects (Chen et al., 2021; Jiang et al., 2024). More recently, the field has seen a shift from human-centric code completion to agentic workflows in which LLMs plan, act, and call tools to solve more realistic software engineering tasks (Liu et al., 2025). These agentic systems—now widely available in production tools (Shrivastava, 2025)—can complete end-to-end workflows that previously required a human developer. Notably, these workflows include raising pull requests (PRs), allowing agents to propose more substantial code changes than earlier bots (He et al., 2023). Producing such substantial changes requires agents to engage with the same software ecosystems and abstractions relied upon by human developers.
@@
Libraries are central to modern software engineering (Somerville, 2016), where reusing established components improves developer productivity and avoids reinventing well-tested functionality (Xu et al., 2020). For agents to contribute meaningfully to real-world projects, they must therefore be able to use libraries effectively. However, prior work shows that LLMs in non-agentic settings often struggle in this respect: exhibiting limited diversity in their library choices (Twist et al., 2025a), omitting version constraints (Raj and Costa, 2024), or calling deprecated APIs (Wang et al., 2025a). To mitigate these issues, existing approaches frequently rely on retrieval-based mechanisms (Tao et al., 2025; Wang et al., 2025b). Agentic systems, however, have greater autonomy and may bypass such additional infrastructure. What remains unclear is whether agentic workflows will actually use external libraries—and if they do, whether this autonomy helps mitigate existing problems or simply reproduces them at scale.
@@
To begin to fill this gap, we present the first focused empirical study of how AI coding agents use libraries in practice, by mining real agent-authored PRs from the AIDev dataset (Li et al., 2025). Specifically, we analyse file diffs (both code and dependency manifests) from 26,760 PRs across four languages (TypeScript, Python, Go, and C#), to answer three research questions (RQs):
@@
RQ1 Library Usage: How often do agents import libraries when generating code, and do they favour built-in or external libraries?
@@
RQ2 New Dependencies: How often do agents introduce new dependencies into a repository, and do they specify versions?
@@
RQ3 Choosing Libraries: Which specific libraries do agents choose, and what functional roles do they play?
@@
Our findings are concise but instructive. 1) Agents often import libraries (29.5% of PRs) but rarely add entirely new dependencies (1.3% of PRs). 2) Agentic workflows encourage far better version hygiene than traditional LLM prompting: agents specify versions for 75.0% of added dependencies, whereas LLMs would only mention versions in $\approx$9% of library-related interactions (Raj and Costa, 2024). 3) Agents use a diverse range of libraries, importing 3,988 distinct external packages across 7,888 PRs—a sharp contrast to the narrow “library preferences” previously observed in more self-contained LLM tasks (Twist et al., 2025a).
Our contributions are as follows.
@@
Language-specific details used for data extraction, including code file extensions, dependency manifest files, and representative import and dependency syntax for TypeScript, Python, Go, and C#.
Programming Language Code File Extensions Example Import Statements Dependency Manifests Example Dependencies
TypeScript (TypeScript, 2025) .ts, .tsx, .js, .jsx, .mjs, .cjs import { useEffect } from ’react’; import { Class } from ’@angular/core’; package.json "lodash": $\sim$4.17.21",
Python (Python, 2025) .py import numpy as np from django.db import models requirements.txt requests pandas<2.0
pyproject.toml "pytest" "numpy==2.79",
Go (Go, 2025) .go import "fmt" import ( mux "github.com/gorilla/mux" ) go.mod require github.com/pkg/errors v0.9.1
C# (Microsoft, 2025) .cs using Newtonsoft.Json; using Serilog; {name}.csproj packages.config <PackageReference     Include="Serilog" Version="2.12.0" />
paket.dependencies nuget Newtonsoft.Json 13.0.1
@@ 2. Dataset
To conduct our study, we use the AIDev dataset (Li et al., 2025), a large-scale corpus of almost a million pull requests (PRs) from over 116,000 GitHub repositories, authored by five distinct LLM-based coding agents (Claude Code, Cursor, Devin, Copilot, OpenAI Codex). Specifically, we use the November 2025 snapshot (commit eee0408 (hao-li, 2025)).
For our analysis, we use the AIDev-Pop subset because it includes the full commit details per PR; AIDev-Pop contains 33,596 PRs from 2,807 high-quality repositories (those with over 100 GitHub stars). We use three tables: repository, containing metadata about each repository; pull_request, which provides high-level PR information; and pr_commit_details, which contains file-level diffs and commit content. The dataset provides diffs in the “combined diff format” (Git, 2025), allowing us to extract new content from each file by selecting lines that begin with whitespace followed by exactly one “+”.
We restrict our study to the four most common languages in the dataset—those with the largest number of repositories—to ensure both depth of analysis and sufficient statistical support. The chosen languages are TypeScript/JavaScript (combined due to nearly identical syntax and library ecosystems, referred to as TypeScript throughout the paper), Python, Go and C#. We initially included Rust (fifth most popular language), but its PRs relied heavily on internal workspace crates that were indistinguishable from external imports, so we limited our study to the top four languages.
Ethical considerations.
@@
Summary statistics for 26,760 agent-authored pull requests, showing per-language counts and percentages for library imports, standard versus external library usage, new dependency additions, version specification, and pull request merge rates.
Metric TypeScript Python Go C# All languages
General statistics Total repositories 840 530 242 220 1,832
Total pull requests 7,480 7,190 10,107 1,983 26,760
Total PRs merged 4,428 (59.2%) 5,468 (76.1%) 8,361 (82.7%) 1,146 (57.8%) 19,403 (72.5%)
RQ1 Library Usage PRs importing any library 2,993 (40.0%) 2,720 (37.8%) 1,273 (12.6%) 902 (45.5%) 7,888 (29.5%)
PRs importing standard library 828 (11.1%) 2,383 (33.1%) 1,244 (12.3%) 743 (37.5%) 5,198 (19.4%)
PRs importing external libraries 2,847 (38.1%) 1,914 (26.6%) 324 (3.2%) 738 (37.2%) 5,823 (21.8%)
Avg. standard library imports per PR 0.26 1.28 0.63 0.49 0.69
Avg. external library imports per PR 1.87 0.73 0.08 0.73 0.80
Unique external libraries imported 2,495 948 314 231 3,988
PRs with library imports merged 1,595 (53.3%) 1,768 (65.0%) 918 (72.1%) 467 (51.8%) 4,748 (60.2%)
RQ2 New Dependencies PRs adding new dependencies 9 (0.1%) 200 (2.8%) 16 (0.2%) 120 (6.1%) 345 (1.3%)
Avg. new dependencies per PR 0.01 0.17 0.02 0.30 0.08
Total new dependencies 95 1,237 203 593 2,128
Dependencies specifying version 95 (100.0%) 1,020 (82.5%) 203 (100.0%) 278 (46.9%) 1,596 (75.0%)
Unique new dependencies 85 552 188 294 1,119
PRs with new dependencies merged 6 (66.7%) 131 (65.5%) 5 (31.2%) 56 (46.7%) 198 (57.4%)
@@ 3.1. Method for RQ1 Library Usage
We begin by investigating how inclined agents are to make use of libraries in their code. We recognise library usage only when imports appear in commits, because imports that already existed in a file prior to the agent’s PR do not reflect the agent’s preferences; and with only partial file snapshots provided in diffs, downstream function calls are often ambiguous without the corresponding import statement. We use file extensions to identify code files for each programming language, as defined by GitHub Linguist (github-linguist, 2025).
We implement language-specific regex parsers to automatically extract import statements from those files, taking care to distinguish relative imports (from within the project itself). All extraction is supported by unit tests to ensure reliability. Information on the file extensions we consider, with example imports, is provided in Table 1. Each library is categorised as either standard (built-in, defined in the language documentation) or external (installed via a third-party package manager), allowing us to measure how much agents rely on external ecosystems versus native language capabilities.
@@ 3.2. Method for RQ2 New Dependencies
Next, we examine how often agents introduce new dependencies into repositories. Here, we are particularly interested in the quality of these additions–specifically, whether agents are explicit with version constraints–because prior work shows that LLMs in non-agentic workflows rarely discuss versions (Raj and Costa, 2024), despite the known risks. For each language, we identify its corresponding dependency manifest files (for example, “requirements.txt” for Python, “package.json” for TypeScript) following the GitHub Docs (GitHub, 2025); we ignore lock files, which contain large numbers of second-hand dependencies not explicitly used within the repository.
We implement manifest-specific regex parsers to automatically identify new dependencies–including any explicit version specifiers (for example, “==1.2” or “¿=1.0”)–again with unit tests to ensure reliability. The manifest files we consider, with example dependencies, are listed in Table 1. This enables us to quantify how often agents introduce new dependencies and how reliably they do so.
@@ 3.3. Method for RQ3 Choosing Libraries
Finally, we examine the specific libraries agents choose and the functional roles they play. For each language, we take the set of libraries identified in RQ1 and RQ2 and compute simple frequency statistics to surface the most commonly used or newly introduced libraries. We then perform manual analysis on the top ten libraries for each language and RQ to determine the areas that agents are most likely to use external libraries.
@@ 4. Findings & Discussion
We analyse a total of 26,760 agent-authored pull requests across 1,832 repositories. Here, we present and discuss our findings.
@@ 4.1. Findings for RQ1 Library Usage
For this RQ, we extract imports from agent-authored code to investigate their library usage patterns. Results in Table 2.
Agents often import libraries when generating code, with 29.5% of PRs importing at least one library, and an average of 1.5 libraries imported across all PRs. However, the extent varies substantially by language: agents lean most heavily on libraries when working in TypeScript and C# (40.0% and 45.5% of PRs), while Go is a clear outlier, importing libraries in just 12.6% of PRs. This suggests that agents primarily perform localised modifications in Go, potentially reflecting both the language’s strong standard library and ecosystem norms that discourage heavy external dependencies. This pattern appears again when considering average imports per PR: both TypeScript and Python exceed two libraries per PR, whereas Go averages below one. Although reusing libraries is a core skill for professional developers, there is no established baseline for how often this should occur, making it difficult to assess whether these rates indicate overuse, underuse, or neither. Despite these substantial cross-language differences, PRs that import libraries are merged at rates only 6–11% lower than PRs that do not, suggesting that library usage itself is not strongly penalised during review.
Agents rely strongly on the standard library in Python and C# (33.1% and 37.5% of PRs), while their use of external libraries in those languages remains comparatively balanced (26.6% and 37.2%). In TypeScript, the pattern is very different: agents overwhelmingly favour external packages (38.1% of PRs) over the standard library (11.1%), indicating a strong preference for high-level tooling. In contrast, Go uses standard libraries almost four times more often than external libraries. Across all languages, agents collectively import 3,988 unique external libraries, indicating substantial diversity. This contrasts sharply with work showing that LLMs typically exhibit low library diversity (Twist et al., 2025a), suggesting that rich project-level context and greater autonomy encourages broader library use. While it may seem surprising that TypeScript alone contributes more than half of all unique libraries, this aligns with the ecosystem’s well-known fragmentation and reliance on micro-packages (Kula et al., 2017); however, such breadth may also increase the risk of inconsistent dependency choices across PRs.
RQ1 Summary. Agents often import libraries for TypeScript (40.0% of PRs), Python (37.8%) and C# (45.5%), but rarely for Go (12.6%). For TypeScript, agents favour external libraries over built-in functionality; for Python and C#, library usage is more balanced; and for Go, the standard library is strongly preferred.
@@
Lists of the ten most frequently imported external libraries and the ten most frequently added new dependencies for each language, along with the number of pull requests in which each library appears.
Metric TypeScript Python Go C#
Top 10 external library imports react (1013) vitest (483) zod (313) next (303) lucide-react (261) @testing-library/react (136) @tanstack/react-query (111) lodash (107) drizzle-orm (97) chai (92) pytest (812) pydantic (279) crewai (264) requests (125) numpy (125) fastapi (114) openai (65) pandas (60) httpx (53) yaml (52) github.com/stretchr/testify (108) github.com/smacker/go-tree-sitter (21) github.com/spf13/cobra (20) github.com/cometbft/cometbft (13) github.com/microsoft/typescript-go (13) github.com/tliron/glsp (13) go.uber.org/zap (12) github.com/gofiber/fiber (11) k8s.io/apimachinery/pkg (10) github.com/onsi/gomega (9) Avalonia (275) Xunit (243) Dock (131) Aspire (55) Azure (50) NUnit (48) FluentAssertions (30) ReactiveUI (28) Newtonsoft (27) Svg (27)
Top 10 new dependencies typescript (3) vite (3) husky (2) @biomejs/biome (2) zod (2) prettier (2) @types/node (2) three (2) @electron-forge/plugin-auto-unpack-natives (1) eslint-plugin-react-compiler (1) requests (24) openai-agents (22) openai (21) numpy (21) pandas (18) pytest (14) python-dotenv (13) uvicorn (13) pydantic (12) aiohttp (12) github.com/golang/snappy (6) github.com/stretchr/testify (3) gopkg.in/yaml.v3 (3) github.com/klauspost/reedsolomon (3) github.com/gogo/protobuf (3) go.uber.org/atomic (2) github.com/smacker/go-tree-sitter (2) github.com/leonklingele/grouper (1) github.com/sashamelentyev/usestdlibvars (1) github.com/breml/bidichk (1) Microsoft.NET.Test.Sdk (26) xunit.runner.visualstudio (22) xunit (20) Moq (20) coverlet.collector (15) Microsoft.AspNetCore.Mvc.Testing (11) NUnit (9) NUnit3TestAdapter (9) Azure.Identity (7) Microsoft.Extensions.DependencyInjection (7)